OpenStudy (torikb232):

What is the value of the expression? (–4)4

OpenStudy (torikb232):

@TheSmartOne

TheSmartOne (thesmartone):

A negative number times a positive number will be negative what is 4 * 4 = ? And then just a negative to it Basically -4 * 4 = -1 * 4 * 4 = ??

OpenStudy (torikb232):

-16?